### v1

# Tree-Planted Transformers: Large Language Models with Implicit Syntactic Supervision

Abstract
Large Language Models (LLMs) have achieved remarkable success thanks to scalability on large text corpora, but have some drawback in training efficiency. In contrast, Syntactic Language Models (SLMs) can be trained efficiently to reach relatively high performance thanks to syntactic supervision, but have trouble with scalability. Thus, given these complementary advantages of LLMs and SLMs, it is necessary to develop an architecture that integrates the scalability of LLMs with the training efficiency of SLMs, namely Syntactic Large Language Models (SLLM). In this paper, we propose a novel method dubbed tree-planting: implicitly “plant” trees into attention weights of Transformer LMs to reflect syntactic structures of natural language. Specifically, Transformer LMs trained with tree-planting will be called Tree-Planted Transformers (TPT), which learn syntax on small treebanks via tree-planting and then scale on large text corpora via continual learning with syntactic scaffolding. Targeted syntactic evaluations on the SyntaxGym benchmark demonstrated that TPTs, despite the lack of explicit syntactic supervision, significantly outperformed various SLMs with explicit syntactic supervision that generate hundreds of syntactic structures in parallel, suggesting that tree-planting and TPTs are the promising foundation for SLLMs.
Tree-Planted Transformers:
Large Language Models with Implicit Syntactic Supervision
Ryo Yoshida and Taiga Someya and Yohei Oseki The University of Tokyo {yoshiryo0617, taiga98-0809, oseki}@g.ecc.u-tokyo.ac.jp
1 Introduction
Recent years have witnessed remarkable success in Large Language Models (LLMs) based on Transformer LMs (Vaswani et al., 2017). The success of LLMs suggests that continual learning on large text corpora is essential for LMs to acquire a wide range of world knowledge and solve various downstream tasks. However, despite their success, LLMs have some drawback in training efficiency. For example, GPT-3 (Brown et al., 2020) is trained on around larger data than a 12-year-old human would have experienced (Warstadt et al., 2023), indicating that LLMs require tremendous training corpus and computational resources.
On another strand, previous work has revealed that Syntactic Language Models (SLMs), the integration of LMs with explicit syntactic supervision, can achieve high performance under data-constrained settings (Dyer et al., 2016; Noji and Oseki, 2021; Qian et al., 2021; Sartran et al., 2022; Yoshida and Oseki, 2022; Murty et al., 2023). For example, Sartran et al. (2022) showed that some SLMs can achieve comparable syntactic knowledge to an LLM-like model111Due to the rapid advances in recent years, what were once considered LLMs are no longer deemed “large” by current standards. We will refer to Transformer LMs larger than or equal to GPT-2 (Radford et al., 2018) as LLM-like models. that is trained with medium—around larger—data, suggesting that syntactic supervision is essential for LMs to achieve high training efficiency. However, despite their training efficiency, SLMs have trouble with scalability: small SLMs cannot compete with LLMs trained on larger data. Thus, given these complementary advantages of LLMs and SLMs, it is necessary to develop an architecture that integrates the scalability of LLMs with the training efficiency of SLMs, namely Syntactic Large Language Models (SLLM; Table 1).
In this paper, we propose a novel method dubbed tree-planting:222The term “tree-planting” coincidentally bears a resemblance to the term used in Mueller and Linzen (2023), but this work diverges from ours in its motivation. Specifically, Mueller and Linzen (2023) investigated biases that enable syntactic generalization in Transformer LMs, from the perspectives of architectural features (depth, width, and number of parameters), as well as the genre and size of training corpus. implicitly “plant” trees into attention weights of Transformer LMs to reflect syntactic structures of natural language. Specifically, Transformer LMs trained with tree-planting will be called Tree-Planted Transformers (TPT), which learn syntax on small treebanks via tree-planting and then scale on large text corpora via continual learning with syntactic scaffolding (Figure 1). Targeted syntactic evaluations on the SyntaxGym benchmark demonstrated that TPTs, despite the lack of explicit syntactic supervision, significantly outperformed various SLMs with explicit syntactic supervision that generate hundreds of syntactic structures in parallel, suggesting that tree-planting and TPTs are the promising foundation for SLLMs.
2 Background
2.1 Large Language Model
Large Language Models are typically based on Transformer LMs (Vaswani et al., 2017) with a large number of parameters and trained on vast amounts of data. A major reason that Transformer LMs are employed as the base architecture for LLMs is their self-attention mechanism, which enables efficient parallel computation on GPUs.
The self-attention mechanism of Transformer LMs computes a representation for predicting the next token through a weighted sum of each token in the context. Specifically, when predicting the -th token, the attention weights from the -th token to the -th token is computed as follows:
where and represent the query vector of the -th token and the key vector of the -th token, respectively, and denotes the dimension of the key vector. As Equation 1 shows, the computation for the -th token prediction does not depend on any computation for the -th token predictions, which enables efficient parallel computation. This property of the self-attention mechanism enables the development of LLMs but it is important to note that these models do not employ any syntactic supervision, although syntactic structures are one of the fundamental properties of natural languages Chomsky (1957).
| Scalability | Training efficiency | |
| LLM | ✓ | |
| SLM | ✓ | |
| SLLM | ✓ | ✓ |
| Parser-free inference | Syntactic supervision | Unidirectional LM | Parallel computation | |
| Wu et al. (2018);Nguyen et al. (2020); Bugliarello and Okazaki (2020);Bai et al. (2021); Sachan et al. (2021);Slobodkin et al. (2022) | ✓ | ✓ | ||
| Wang et al. (2019) | ✓ | ✓ | ||
| Strubell et al. (2018);Chen et al. (2023) | ✓ | ✓ | ✓ | |
| Peng et al. (2019) | ✓ | ✓ | ✓ | |
| Tree-planting (ours) | ✓ | ✓ | ✓ | ✓ |
2.2 Syntactic Language Model
Syntactic Language Models (SLMs) are a generative model of a token sequence and its syntactic structure :
where denotes the sequence of actions to generate both the token sequence and syntactic structure. For example, in top-down and left-to-right SLMs, each could be either generating a token, opening a constituent, or closing a constituent.
Recently, several SLMs based on the Transformer architecture have been proposed, achieving higher syntactic knowledge than medium LLM-like models Qian et al. (2021); Sartran et al. (2022); Murty et al. (2023). However, because SLMs model the joint probability of a token sequence and its syntactic structure, they cannot be trained on other than treebanks, which prevents them from scaling on large text corpora Moreover, SLMs also have practical drawback in inference costs: when utilized as LMs, they require hundreds of syntactic structures via beam search (Stern et al., 2017; Crabbé et al., 2019) or an external parser, to marginalize joint distribution to precisely approximate .
2.3 Constraints on attention weights
As discussed in Subsection 2.2, the bottleneck that prevents SLMs from scaling on large text corpora is their modeling space of the joint probability. To achieve the foundational architecture for SLLMs, it is necessary to introduce syntactic knowledge without changing the modeling space of their underlying Transformer LMs. For our goal, we will build upon another line of approach that constrains attention weights according to syntactic structures—typically targeting bidirectional Transformer Encoders like BERT (Devlin et al., 2019)—and extend it to unidirectional Transformer LMs. Table 2 summarizes the previous work in this line of approach, comparing our tree-planting (Section 3) against others based on the requirements for SLLM: (i) parser-free inference, (ii) syntactic supervision, (iii) unidirectional LM, and (iv) parallel computation.
First, the majority of these approaches are purely motivated to explicitly restrict attention weights with syntactic structures from external parsers, under the assumption that these parsers would be available during inference (Wu et al., 2018; Nguyen et al., 2020; Bugliarello and Okazaki, 2020; Bai et al., 2021; Sachan et al., 2021; Slobodkin et al., 2022). These studies achieved successful performance in their respective downstream tasks, but not only are they all not directly applicable to unidirectional LMs, they also require external parsers during inference, rendering them unsuitable as the foundation for SLLM.
Second, several studies proposed approaches that do not require external parsers during inference. Wang et al. (2019) aimed at an unsupervised approach, where a hierarchical architectural bias widens the range of neighboring tokens eligible to attend from lower to upper layers, though this method is also not aligned with our goal of achieving higher training efficiency via syntactic supervision. Furthermore, Strubell et al. (2018) and Chen et al. (2023) designed the loss functions that implicitly encourage the attention to syntactic parents or children for each token, satisfying the 3/4 requirements for SLLM. However, this approach is potentially not suitable for unidirectional LMs where the existence of the dependent in the left context is not guaranteed.
Finally, the approach also most closely aligned with the spirit of this research is a hybrid Parser and neural Language Model (PaLM; Peng et al., 2019). PaLM is the integration of an RNN LM with an additional attention layer, which would be supervised to attend the constituent spans among the spans ending at time : . Although PaLM also meets the 3/4 requirements, it was by nature proposed for RNN LMs. The challenge arises when adapting PaLM to Transformer LMs; the generation of embeddings for the spans introduces a significant bottleneck in parallel computation with the self-attention mechanism.
To sum up, none of the previous approaches fully satisfy the requirements for SLLM, highlighting the necessity for innovative methodologies.
3 Proposed method: tree-planting
In this paper, we propose a novel method dubbed tree-planting: implicitly “plant” trees into attention weights of Transformer LMs to reflect syntactic structures of natural language (Figure 2). Specifically, Transformer LMs trained with tree-planting will be called Tree-Planted Transformers (TPT), which learn syntax on small treebanks via tree-planting and then scale on large text corpora via continual learning with syntactic scaffolding. Tree-planting is strictly designed to satisfy the requirements for SLLM: (i) parser-free inference, (ii) syntactic supervision, (iii) unidirectional LM, and (iv) parallel computation.
3.1 Supervision of attention weights
In producing the supervision of attention weights, we extend the notion of syntactic distance (Shen et al., 2018, 2019; Du et al., 2020), a 1D sequence of the number of edges on syntactic structures between two consecutive words, to a 2D matrix between all pairs of words:
where and represent the -th and -th words, respectively, and is the function that maps a pair of words to the number of edges on syntactic structures between them. This notion of syntactic distance matrix is theory-neutral: applied to any kind of syntactic structure, as long as the number of edges can be counted on it.333When applied to dependency structures, we ignore the direction of syntactic dependency.
Then, the syntactic distance matrix is converted to the supervision of attention weights as follows:
where represents the supervision of the attention weight from the -th word to the -th word when predicting the -th word. This design of the supervision expects the attention weight of each word to decrease exponentially with its number of edges between the predicted word;444We adopt an exponential function as Lin and Tegmark (2017) reported that the mutual information between words will decay exponentially with respect to the number of edges on the syntactic structure between them. this alone successfully satisfies the 3/4 requirements for SLLM: (ii) syntactic supervision, (iii) unidirectional LM, and (iv) parallel computation. To fulfill the remaining requirement of (i) parser-free inference, we adopt a strategy similar to that of Strubell et al. (2018); Chen et al. (2023), designing the loss function to implicitly supervise attention.
3.2 Loss function
The supervision in Subsection 3.1 is produced at the word level but LLMs typically take their input at the subword level. To bridge this gap, we first convert the subword-level attention weight matrix from a targeted Transformer LM to the word-level attention weight matrix as follows:
where represents the word-level attention weight from the -th word to the -th word. is defined as the sum of the subword-level attention weights over the subword inside when predicting the subword inside , with representing the subword-level attention weight from the -th subword to the -th subword and and being the functions that map words to their start and end subword index, respectively. We employ from specific attention heads called tree-planted heads.555Qian et al. (2021) also proposed the architecture which constrains some attention heads based on syntactic structures, or PLM-mask. PLM-mask and our tree-planting are similar in spirit, but they are quite different in their implementation: PLM-mask is a type of SLM that jointly generates a word sequence and its syntactic structure, but tree-planting builds TPTs, a type of LM. Furthermore, PLM-mask explicitly masks the attention weights based on the local parser state but tree-planting implicitly guides attention weights to reflect the whole syntactic structure.
To implicitly supervise the word-level attention weight matrix with the supervision , we introduce a tree-planting loss employing a Kullback–Leibler (KL) Divergence loss :666This loss function is inspired by Ma et al. (2023), which guides attention weights to focus on relevant texts in a document-level relation extraction task.
where represents the length of a word sequence . In short, the tree-planting loss is the average KL Divergence loss in predicting each word except the beginning of .
During the training, is averaged over tree-planted heads and balanced with the next word prediction loss :
where represents a tree-planting loss for each tree-planted head , is the total number of tree-planted heads, and is a weight that balances the importance of the next word prediction loss and the average tree-planting loss. Transformer LMs trained with this loss function will be called Tree-Planted Transformers (TPT).
4 Experiment
To confirm that syntactic knowledge is introduced to TPTs by tree-planting, we conduct training on a small treebank and targeted syntactic evaluations on a syntactic knowledge benchmark.
4.1 Settings
Training data
We used LG dataset of Hu et al. (2020), which comprises approximately 42M tokens from BLLIP corpus (Charniak et al., 2000). Implicit syntactic supervision with each of three types of syntactic structures was investigated: (i) dependency structures ([dep.]), (ii) constituency structures ([cons.]), and (iii) binarized constituency structures ([bin.]). The (i) dependency structures were parsed with the en_core_web_sm model from the spacy library (Montani et al., 2023).777https://spacy.io The (ii) constituency structures were re-parsed with the Berkeley Neural Parser (Kitaev and Klein, 2018)888https://github.com/nikitakit/self-attentive-parser by Hu et al. (2020). The (iii) binarized constituency structures were obtained by the binarization of the (ii) constituency structures with the chomsky_normal_form function from the nltk library (Bird et al., 2009).999https://www.nltk.org
Models
We used the same architecture and BPE tokenizer as GPT-2 small (124M; Radford et al., 2018). The implementation of GPT2LMHeadModel and GPT2Tokenizer from the transformers library (Wolf et al., 2020)101010https://huggingface.co/docs/transformers were employed but all parameters of GPT2LMHeadModel were randomly initialized. For the tree-planted head and the weight of the tree-planting loss, we adopted a single attention head on the last layer and , respectively. The choice of the tree-planted head and the weight was based on preliminary experiments and the detailed effects of them will be described in Section 5.
As baselines, we trained three models: (i) a model with zero weight for the tree-planting loss ([zero]), (ii) a model supervised with random syntactic distances that were generated from the distribution same as the dependency structures ([rand.]), and (iii) a model supervised with sequential distances ([seq.]). Note importantly, (i) is equivalent to a Transformer LM. Hyperparameters are shown in Appendix A.
Evaluation data
[…]
The author next to the senators is good.
[…]
All models were trained and evaluated two times with different random seeds. We report average accuracies with a standard deviation, along with word-level perplexity on the BLLIP test set.
4.2 Overall accuracies
| SG () | PPL () | |
| Baselines: | ||
| TPT[zero] | 71.7 0.3 | 47.5 0.1 |
| TPT[rand.] | 69.0 1.0 | 47.4 0.1 |
| TPT[seq.] | 70.1 3.5 | 47.3 0.2 |
| TPTs (ours): | ||
| TPT[dep.] | 77.1 0.2 | 47.7 0.1 |
| TPT[cons.] | 75.8 0.0 | 45.5 0.0 |
| TPT[bin.] | 73.0 1.8 | 45.6 0.2 |
| SLMs (comparable): | ||
| PLM | 42.2 1.2 | - |
| PLM-mask | 42.5 1.5 | - |
| SLMs (reference): | ||
| PLM† | 73.2 0.6 | 49.3 0.3 |
| PLM-mask† | 74.6 1.0 | 49.1 0.3 |
| TG‡ | 82.5 1.6 | 30.3 0.5 |
| LLM-like models (reference): | ||
| GPT-2¶ | 78.4 | - |
| Gopher¶ | 79.5 | - |
| Chinchilla¶ | 79.7 | - |
Table 3 shows the overall accuracies of TPTs and their baselines on the SyntaxGym benchmark (SG), along with word-level perplexity on the BLLIP test set (PPL). The overall accuracies were calculated across the syntactic suites. We also report the accuracies of several SLMs that were also trained on the same BLLIP-LG dataset: PLM, PLM-mask (Qian et al., 2021), and TG (Sartran et al., 2022). Only unmarked PLM and PLM-mask can be fairly comparable with TPTs as their evaluation was conducted generating a single syntactic structure via greedy search, to align inference costs with TPTs.111111The fair comparison of TG was not performed because their trained parameters were not publicly available. and represent the reference points from Sartran et al. (2022) as their inference methods are more costly than TPTs: and employed word-synchronous beam search (Stern et al., 2017) of action beam size 100121212Word beam size was 10 and fast track size was 5. and the external parser (Dyer et al., 2016), respectively. The accuracies of several LLM-like models are also reported from Sartran et al. (2022): GPT-2 (Radford et al., 2018), Gopher (Rae et al., 2022), and Chinchilla (Hoffmann et al., 2022). They are also the reference points as these LLM-like models were trained on significantly larger corpora (denoted by ). Perplexity can be directly comparable only within the same mark, either or , due to differences in the tokenization of the constituency parser and dependency parser.
There are some important observations in the overall accuracies on the SyntaxGym benchmark:
TPT[zero], which is equivalent to a Transformer LM, underperformed all TPTs with some implicit syntactic supervision, suggesting that tree-planting can introduce syntactic knowledge to TPTs.
TPTs[rand.][seq.] also underperformed all TPTs with some implicit syntactic supervision, indicating that not KL Divergence loss itself but the loss based on syntactic structures is necessary.
Among TPTs with some implicit syntactic supervision, TPT[dep.] achieved the best performance. We further investigate this point in Subsection 4.3.
Most importantly, despite the lack of explicit syntactic supervision, TPTs[dep.][cons.] significantly outperformed not only the comparable SLMs (unmarked PLM and PLM-mask) but also the various SLMs that generate hundreds of syntactic structures in parallel (PLM and PLM-mask).
Even though the best TPT[dep.] underperformed the reference points of the more costly TG and the larger LLM-like models, these observations adequately suggest that tree-planting and TPTs are the promising foundation for SLLMs.
Regarding perplexity, although TPT[dep.] numerically underperformed its comparable baselines, they all achieved similar perplexity with no significant differences.
4.3 Circuit accuracies
In this subsection, we investigate the advantages of dependency structures through the lens of circuit accuracies. Figure 3 shows the circuit accuracies of TPTs with some implicit syntactic supervision and the baseline model with zero weight for the tree-planting loss on the SyntaxGym benchmark. The circuit accuracies calculated across the syntactic suites (the vertical axis) are plotted against the models (the horizontal axis), with each dot denoting the accuracy of a specific seed.
vs. zero supervision
TPT[dep.] outperformed TPT[zero] on 5/6 circuits, suggesting that syntactic supervision of dependency structures is generally advantageous over zero supervision. However, the Garden-Path Effects circuit presents an exception, where LMs are evaluated for the ability to be surprised in a human-like manner, through comparisons between sentences minimally different not in grammaticality but in local ambiguity (Hu et al., 2020). The underperformance of TPT[dep.] may suggest that due to the syntactic knowledge introduced by tree-planting with dependency structures, TPT[dep.] was no longer surprised by locally ambiguous but grammatical sentences. We further investigate this point in Appendix B.
vs. constituency structures
[…]
vs. binarized constituency structures
TPT[dep.] outperformed TPT[bin.] on 3/6 circuits, with similar performance (a difference less than ) on the other 3 circuits. Notably, TPT[dep.] achieved significantly better performance (a difference more than ) on the Agreement and Licensing circuits. Noji and Oseki (2021) reported that deep syntactic supervision is not always optimal; rather mild syntactic supervision is sufficient for addressing long-distance dependencies between elements within and outside complex NP subjects. Given that (i) the Agreement and Licensing circuits consist only of syntactic suites that exemplify this condition131313Among the other syntactic circuits, the Center Embedding circuit also exemplifies this condition. and (ii) the average syntactic distance in the training data is significantly shorter for dependency structures (4.8) than binarized constituency structures (13.1), it could be argued that dependency structures would be more suitable as “good enough” syntactic supervision than binarized constituency structures.141414The average syntactic distance of constituency structures is 10.0. This suggests that dependency structure would also be superior to constituency structure as “good enough” syntactic supervision, besides the points discussed in the “vs. constituency structures” paragraph.
5 Analysis
In this section, we report the effects of (i) the number of tree-planted heads and (ii) the weight of a tree-planting loss, using TPT[dep.].
5.1 Number of tree-planted heads
Our TPTs are based on a 12-layer, 12-head Transformer LM. In Section 4, out of heads, we adopted a single attention head on the last layer as a tree-planted head. In this subsection, we explore two alternatives: (i) head-direction extension and (ii) layer-direction extension. For the head-direction extension, 0, 1, 3, 6, 9, and 12 heads on the last layer were adopted as tree-planted heads. For the layer-direction extension, one attention head from the each of bottom 0, 1, 3, 6, 9, and 12 layers was adopted as tree-planted heads.
In the left two columns of Figure 4, the results of the head-direction and layer-direction extension are shown: the overall accuracies on the SyntaxGym benchmark (SG) and the word-level perplexity on the BLLIP test set (PPL) (the vertical axis) are plotted against the number of tree-planted heads (the horizontal axis). Each dot denotes the accuracy or perplexity of a specific seed. For both settings, are equivalent to TPT[zero] and TPT[dep.], respectively.
Considering the overall accuracies on the SyntaxGym benchmark, in both the head-direction and layer-direction extension, the highest accuracy was achieved when only a single head was adopted as a tree-planted head, while it is noteworthy that all the models with tree-planted heads outperformed the model without them. Incidentally, it should be mentioned that the result of the layer-direction extension exhibited significantly more variability. Although the reason why a single tree-planted head would work well is unclear, the adoption of multi tree-planted heads inherently induces the handling of redundant information across heads, which might potentially hinder the management of non-syntactic information of natural languages (e.g., lexical information). Regarding perplexity, no consistent trend emerged.
5.2 Weight of a tree-planting loss
In Section 4, we adopted as the weight of the tree-planting loss. Here, we extend to , , , , and . are equivalent to TPT[zero] and TPT[dep.], respectively.
The rightmost column of Figure 4 shows the results of the weight extension. The overall accuracies on the SyntaxGym benchmark display a single-peaked pattern, with the maximum reached for . Interestingly, this result suggests that by overtly focusing on reflecting syntactic structures, TPTs paradoxically become unable to learn syntactic knowledge. On the other hand, we observed that the perplexity got worse monotonically as the weight increased. From these observations, we may deduce that to acquire syntactic knowledge, TPTs should learn not only to reflect syntactic structures in their attention weights but also to precisely predict the next word. Therefore, the weight of the tree-planting loss emerges as a critical hyperparameter, indicating that the search for the optimal balance between the next-word prediction loss and tree-planting loss is vital for developing more human-like TPTs.
6 Conclusion
In this paper, we proposed a novel method dubbed tree-planting: implicitly “plant” trees into attention weights of Transformer LMs to reflect syntactic structures of natural language. Specifically, Transformer LMs trained with tree-planting are called Tree-Planted Transformers (TPT), which learn syntax on small treebanks via tree-planting and then scale on large text corpora via continual learning with syntactic scaffolding. Targeted syntactic evaluations on the SyntaxGym benchmark demonstrated that TPTs, despite the lack of explicit syntactic supervision, significantly outperformed various SLMs with explicit syntactic supervision that generate hundreds of syntactic structures in parallel, suggesting that tree-planting and TPTs are the promising foundation for SLLMs.
Limitations
There are at least three limitations in this paper. First, we only conducted sentence-level tree-planting. Typically, LLMs are trained at the document level, but SLMs are trained at the sentence level (Dyer et al., 2016; Kuncoro et al., 2017; Noji and Oseki, 2021; Yoshida and Oseki, 2022), because on treebanks, the annotations are assigned at the sentence level. Because of this constraint, we also employed sentence-level experimental design and verified the effectiveness of the proposed method first and foremost. Recent research in SLMs, however, has begun to extend treebank annotations to the document level and train document-level SLMs on them (Sartran et al., 2022; Murty et al., 2023). When constructing TPTs for practical use, it might be beneficial to follow these recent studies and perform tree-planting with document-level annotations.
Second, we only evaluated TPTs on the syntactic knowledge benchmark and perplexity. Recently, Murty et al. (2023) evaluated the performance of SLMs on tasks other than the targeted syntactic evaluations for the first time, suggesting that syntactic knowledge could also be beneficial to solving them. This indicates that there is also room for a broader evaluation of our methodology.
Finally, the development of a novel continual learning method (e.g., updating the parameters of tree-planted heads sparingly) would be necessary for scaling TPTs on large corpora, without compromising the syntactic knowledge but rather exploiting it as syntactic scaffolding. In future work, we plan to develop a novel method for "climbing trees" in TPTs.
Ethical considerations
A significant feature of TPT lies in the training efficiency, which can potentially contribute to reducing computational resources. One minor concern is the possibility of bias in the models utilized in this paper, attributed to the training data (i.e., the BLLIP corpus), although this experimental setting follows conventional practices in the literature on SLMs. We employed ChatGPT and Grammarly for writing assistance, and for the development of experimental code, we utilized ChatGPT and Copilot. These tools were used in compliance with the ACL 2023 Policy on the Use of AI Writing Assistance.
Acknowledgements
We would like to thank Peng Qian for providing us with the re-parsed BLLIP-LG dataset, which is used to train TPT[cons.]. We also wish to thank Laurent Sartran for answering various questions. We are also grateful to Kohei Kajikawa, Shinnosuke Isono, Yushi Sugimito, and Taketeru Yamakoshi for their insightful comments and suggestions. This work was supported by JST PRESTO Grant Number JPMJPR21C2.
Appendix A Hyperparameters
Hyperparameters for our experiments are shown in Table 4, which primarily followed default settings. All models were trained and evaluated on NVIDIA V100 (16GB). The total computational cost for all experiments in this paper amounted to about 1,300 GPU hours.
| Optimizer | AdamW |
| Learning rate | 5e-5 |
| Number of epochs | 10 |
| Dropout rate | 0.1 |
| Batch size | 256 |
Appendix B Further investigation of the Garden-Path Effects circuit
In Subsection 4.3, we suggest the probability that syntactic knowledge introduced by tree-planting with dependency structures may prevent TPT[dep.] from being surprised by locally ambiguous but grammatical sentences. To inspect this, we break down the Garden-Path Effects circuit into the syntactic suites: “main verb / reduced relative clause” (MVRR) and “NP/Z garden-paths” (NP/Z).
[…]
The dog seen on the beach chased
after a bird.
!The dog walked on the beach chased
after a bird.
[…]
Conversely, TPT[cons.][bin.] did not underperform TPT(zero.) on the MVRR suites. This result could be straightforwardly understood, given that on these structures, the participle verb (seen/walked) and the head of the subject NP (dog) are equidistant from the main verb (chased). However, it is worth noting that the determiner of the subject NP (the) also shares this distance, which may not always be a desirable property (cf. Subsection 4.3).
Finally, PLM-mask, the more costly SLM, also underperformed TPT[zero] on the MVRR suites. This suggests that the models with explicit syntactic supervision may also struggle with losing sensitivity to the local ambiguity as PLM[dep.].
[…]
[theme = simple]
{deptext}[column sep=0em]
The & author & next & to & the & senators & is/*are & good.
2 & 1 & 2 & 3 & 5 & 4 & 0 & 1
[…]
for tree=
align=center,
parent anchor=south,
child anchor=north,
l=0.1cm, s=0.1cm, inner sep=1pt, ,
[S, for tree=parent anchor=south, child anchor=north
[NP
[NP [DT [The
7, tier=word]] [NN [author
7, tier=word]]]
[ADVP [JJ [next
7, tier=word]] [PP [IN [to
8, tier=word]]
[NP [DT [the
9, tier=word]] [NNS [senators
9, tier=word]]]
[VP [VBZ [is/*are
0, tier=word]] [ADJP [JJ [good.
5, tier=word]]]]
Appendix D Begin/End Of Sentence Tokens
Sentences in the BLLIP corpus do not include Begin/End of Sentence (BOS/EOS) tokens, which are essential for sequences processed by LMs. To integrate these tokens, we implemented the following modifications:
For dependency structures, we introduced BOS/EOS tokens by defining new edges from the ROOT to these tokens.
For constituency structures, we introduced the BOS/EOS tokens by modifying the tree structure to encapsulate the original structure within a new root node, specifically by adding a BOS token and an EOS token as the first and the last child of this new root, respectively.
Appendix E License of the data/tools
We summarize the license of the data/tools employed in this paper in Table 5. All data and tools were used under their respective license terms.
Data/tool |
License |
BLLIP (Charniak et al., 2000) |
BLLIP 1987-89 WSJ Corpus Release 1 License Agreement |
SyntaxGym (Gauthier et al., 2020) |
MIT |
spacy (Montani et al., 2023) |
MIT |
nltk (Bird et al., 2009) |
Apache 2.0 |
transformers (Wolf et al., 2020) |
Apache 2.0 |
Berkeley Neural Parser (Kitaev and Klein, 2018) |
MIT |
PLM/PLM-mask (Qian et al., 2021) |
MIT |